OpenStudy (anonymous):

at 4:00 Gigi and Model left the music house Gigi walked east and reached her house at 5:00 Madel rode her bike west and reached her house at 4:15.madel traveled twice as fast as gigi.their houses are 9km apart.how part did each travel?

OpenStudy (anonymous):

Distance = time*speed Gigi Time = 60 Madel Time = 15 Gigi Speed = x Madel speed = 2x Gigi Distance = time*speed = 60*x = 60x Madel Distance = time*speed = 15*2x = 30x 60x + 30x = 9 90x = 9 x = 1/10 Gigi Distance = 60*1/10 = 6km Madel Distance = 15*2*1/10 = 3km